40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog  Not logged in ELOG logo
Entry  Thu Jun 19 23:16:50 2014, Koji, Optics, General, All of the prisms have been glued P6196525.JPG
    Reply  Fri Jun 20 18:59:12 2014, Koji, Optics, General, All of the invar blocks have been glued 
       Reply  Mon Jun 23 21:54:16 2014, Koji, Optics, General, All of the gluing completed 
Message ID: 187     Entry time: Thu Jun 19 23:16:50 2014     Reply to this: 188
Author: Koji 
Type: Optics 
Category: General 
Subject: All of the prisms have been glued 

- All of the PRISM mirrors have been glued

- 4 out of 5 beam dumps have been glued


TODO

EP30-2 gluing of the INVAR blocks for the PDs

PDs/QPDs need to be slightly lower -> order more shims

Remove the templates

Glue the last beam dump

Vibration test?

Bring the OMC to the bake lab

Vacuum baking

Bring it back to the OMC lab

Cabling / Wiring

Optical tests

Backscattering test

Packing / Shipping

P6196525.JPG

ELOG V3.1.3-